# A subset $S$ of vertices in a planar graph $G$ is a free set if, for every set $P$ of $|S|$ points in the plane, there exists a straight-line crossing-free drawing of $G$ in which vertices of $S$ are mapped to distinct points in $P$. In this survey, we review

### v1

@@ 3.2 Free Sets in Planar Graphs
Therefore, $G_{x}:=G[\{x_{1},\ldots,x_{k}\}]$ is an induced outerplane subgraph of $G$ whose outer face is bounded by the cycle $C_{x}$. By Lemma 4, there is simple closed curve $C$ (which closely follows $C_{x}$) that contains a set $S$ of at least $k/2$ vertices of $G_{x}$, that is contained in the closure of the outer face of $G_{x}$, that contains $G_{x}$ in the closure of its interior, and that is proper and good with respect to $G_{x}$. Since each edge of $G$ with two endpoints in $G_{x}$ is contained in the closure of the interior of $C_{x}$, $C$ (if drawn sufficiently close to $C_{x}$) intersects each edge in $E(G)\setminus E(G_{x})$ in at most one point, as illustrated in Fig. 8(d). Thus $C$ is a proper-good curve for $G$ that contains the vertices in $S$. By 1, $S$ is a free set in $G$ of size at least $k/2$.
@@
Next, consider some maximal antichain $S:=y_{1},\ldots,y_{k}:=v_{i_{1}},\ldots,v_{i_{k}}$ of $(V(G),\prec)$ of length $k>1$ ordered by canonical ordering so that $i_{1}<\cdots<i_{k}$, as illustrated in Fig. 8(e). Since $k>1$, $S$ does not contain $v_{1}$ or $v_{2}$. Since $S$ is maximal and does not contain $v_{1}$ or $v_{2}$, $y_{k}=v_{n}$. Consider the sequence of cycles $C_{1},\ldots,C_{k}$ where $C_{j}$ is the cycle that bounds the outer face of $G_{i_{j}}$. By definition of canonical ordering, the interior of $C_{j}$ contains the interior of $C_{j-1}$ for each $j\in\{2,\ldots,k\}$. The nesting of these cycles is illustrated in Fig. 8(e), where each new colour shows the interior of $C_{j}$ that is not contained in the interior of $C_{j-1}$. Furthermore, the interior of $C_{j}$ must contain the vertex $y_{j-1}$ since, otherwise both $y_{j-1}$ and $y_{j}$ are on the outer face of $G_{i_{j}}$, which would mean that $y_{j-1}$ and $y_{j}$ are comparable. Consider now the union of two cycles $C_{j-1}$ and $C_{j}$ (or rather the union of the two closed curves that represent these two cycles in the crossing-free drawing of $G$). This union has one (bounded) faces, $f$, that contains both $y_{j-1}$ and $y_{j}$ on its boundary. The boundary of $f$ is a cycle $D_{j}$ in $G$ comprised of two paths: a path in $C_{j-1}$ containing $y_{j-1}$ and a path in $C_{j}$ containing $y_{j}$. Note that the union of all the cycles $D_{1},\dots D_{k}$, (or rather the union of the $k$ closed curves that represent these $k$ cycles in the crossing-free drawing of $G$) defines $k+1$ faces in the plane, the weak dual of which is a path. These faces are illustrated in Fig. 8(f), where the interior of each face is assigned its own colour.
@@
We now construct a proper-good curve $C$ for $G$ that contains the vertices of $S$. Let $y_{0}$ be a point in the interior of the edge $v_{1}v_{2}$. Let $G_{1}^{\prime}:=G_{i_{1}}$ and, for each $j\in\{2,\ldots,k\}$, let $G_{j}^{\prime}$ be the near-triangulation whose outer face is bounded by the cycle $D_{j}$ described in the previous paragraph, and whose inner faces are faces of $G$. Graphs $G_{j}^{\prime}$ are illustrated in Fig. 8(f) by having their interiors shaded. Observe that $G^{\prime}_{1},\ldots,G^{\prime}_{k}$ have pairwise disjoint interiors. By Lemma 5, there is a simple open curve $I_{j}$:$(0,1)\to\operatorname{int}(G^{\prime}_{j})$ with endpoints $y_{j-1}$ and $y_{j}$ that is proper with respect to $G^{\prime}_{j}$, for each $j\in\{1,\ldots,k\}$. Since $G^{\prime}_{1},\ldots,G^{\prime}_{k}$ have pairwise disjoint interiors, and since the union of the closures of $\operatorname{int}(G^{\prime}_{1})\ldots,\operatorname{int}(G^{\prime}_{k})$ has a path for its week dual, the open curve obtained by taking the union of the closures of $I_{1},\ldots,I_{k}$ is simple and it is proper with respect to $G$. Finally, let $I_{k+1}$ be a simple open curve with endpoints $y_{0}$ and $y_{k}=v_{n}$ whose interior is contained in the outer face of $G$, as illustrated in Fig. 8(f). Then the simple closed curve $C$ obtained from the union of the closure of open curves $I_{1},\ldots,I_{k+1}$ is a proper-good curve for $G$ that contains $S$. By 1, $S$ is a free set of size $k$.
@@
To complete the proof we use Dilworth’s Theorem, which guarantees that the poset $(V(G),\prec)$ contains a chain of size at least $\sqrt{2n}$ or an antichain of size at least $\sqrt{n/2}$. In either case we obtain a free set of size at least $\sqrt{n/2}$. ∎
@@ 3.3 Free Sets in Max-Degree-$\Delta$ Planar Graphs
Da Lozzo et al. [24] suggest the possibility that, since upper bounds on the circumference of dual graphs can be used to obtain upper bounds on the size of free sets, maybe lower bounds on circumference can be used to prove the existence of large collinear sets. Dujmović and Morin [28] show that, for planar graphs of bounded degree, this is indeed the case. In short, they show that a triangulation $G$ of maximum-degree $\Delta$ and whose dual has circumference $c(G^{*})$ has a free set of size $\Omega(c(G^{*})/\Delta^{4})$. A series of results has steadily improved the lower bounds on the circumference of $n$-vertex (not necessarily planar) cubic triconnected graphs [10, 14, 41, 12, 44]. The current record is held by Liu et al. [44] who show that, for any $n$-vertex cubic triconnected graph $G^{*}$, $c(G^{*})\in\Omega(n^{0.8})$.
In the remainder of this subsection we describe some of the techniques used in [28] to establish the $\Omega(c(G^{*})/\Delta^{4})$ result. Cycles in $G^{*}$ are relevant to free sets because every cycle in $G^{*}$ corresponds to a proper-good curve in $G$. The resulting curve $C$ does not contain any vertices of $G$, but it is natural to try reroute $C$ to obtain a new curve $G$ that goes through some vertices of $G$. Fig. 9 (a) depicts a situation where this rerouting fails because the resulting curve is no longer proper because it intersects the red edge in Fig. 9(b) in two points. This leads to the following definition that lays out conditions under which such rerouting of curve $C$ is safe. We say that a vertex $v$ of $G$ is caressed by $C$ if the edges of $G$ that are incident to $v$ and intersected by $C$ appear consecutively around $v$, as depicted in Fig. 9(c) (the thick grey edges crossed by $C$ are consecutive around $v$ unlike in Fig. 9(a) where the set of thick grey edges crossed by $c$ is not consecutive around $v$). If $C$ caresses $v$ and intersects the edges $vv_{1},\ldots,vv_{r}$ (see Fig. 9(c)) then there is a sequence of faces $f_{0},\ldots,f_{r}$ where $vv_{i}$ is the edge shared between $f_{i-1}$ and $f_{i}$. Then the portion of $C$ that intersects $f_{0},\ldots,f_{r}$ can be replaced with a curve $C^{\prime}$ that enters $f_{0}$, proceeds directly to $v$, and immediate exits $f_{r}$ (see Fig. 9(d)). Because the original curve $C$ does not intersect any other edges incident to $v$, the modified curve $C^{\prime}$ is also a proper-good curve. This operation can be repeated on any set of caressed vertices that form an independent set:
@@
For each $v\in S$, reroute $C$ as described above so that $C$ contains $v$. Since $v$ is caressed by $C$, the rerouting that takes place at $v\in S$ causes $C$ to intersect each edge incident to $v$ in exactly one point, namely $v$, but does not change the intersection of $C$ with any edges not incident to $v$. Since the vertices in $S$ form an independent set, all of these rerouting operations do not cause $C$ to intersect any edge of $G$ in more than one point. Thus $C$ is a proper-good curve that contains $S$ so, by 1 $S$ is a free set in $G$. (The condition that $S$ have at least two vertices avoids the case in which $C$ crosses all the edges incident to a single vertex $v$.) ∎
The requirement that a curve $C$ caresses a vertex $v$ of $G$ is equivalent to requiring that the intersection of $C$ with the face $f_{v}$ of $G^{*}$ that contains $v$ is a path. When this happens, we say that $C$ caresses $f_{v}$. Thus, finding a large collinear set in $G$ is equivalent to finding a cycle in $G^{*}$ that caresses many faces of $G^{*}$.
@@
Let $G$ be a triangulation and let $G^{*}$ be the dual of $G$. If some cycle $C$ in $G^{*}$ caresses $k$ faces of $G^{*}$ then $G$ contains a free set of size at least $k/4$.
@@
Unfortunately, there exists cubic triconnected planar graphs $G^{*}$ with no large faces that have a Hamiltonian cycle that caresses only four faces of $G^{*}$. One such graph is shown in Fig. 10. The main technical result of Dujmović and Morin [28] is to show that such cycles can be modified to produce cycles that caress many faces:
@@
Refer to Fig. 11. The proof of Lemma 8 is far too long to include here in any detail, so we give a high level sketch. For the sake of simplicity, let $\Delta$ be a fixed constant. Let $C$ be a cycle in $G^{*}$. Say that $C$ touches a face of $G^{*}$ if $C$ and $f$ share at least one edge. Since each face of $G^{*}$ has at most $\Delta$ edges and each edge of $C$ touches two faces of $G^{*}$ (one inside $C$ and one outside of $C$), the number of faces of $G^{*}$ that are touched by $C$ is at least $2|C|/\Delta$. At least $|C|/\Delta$ of these faces are in the interior of $C$ and at least $|C|/\Delta$ of these are in the exterior of $C$.
The authors define a subgraph $H$ of $G^{*}$ that includes all the edges of $C$ and such that any cycle in the dual graph $H^{*}$ of $H$ contains faces inside and outside of $C$. Removing the edges from $H^{*}$ that correspond to edges of $C$ produces two trees $T_{0}$ and $T_{1}$, where $T_{0}$ contains faces of $H$ in the interior of $C$ and $T_{1}$ contains faces of $H$ in the exterior of $C$. For each $b\in\{0,1\}$, the tree $T_{b}$ has two important properties:
@@
The second of these properties says that any node of $T_{b}$ that contains many faces touched by $C$ either caresses many of these faces or has high degree in $T_{b}$. The number of leaves in $T_{b}$ is $2+\sum_{x}(\deg_{T_{b}}(x)-2)$, where the sum is over all non-leaf nodes $x$ of $T_{b}$. Combining this with the first of these properties implies that $C$ caresses $\Omega(|C|)$ faces of $G^{*}$ or that $T_{b}$ has $\Omega(|C|)$ nodes but at most $\epsilon|C|$ leaves, for some small $\epsilon>0$. In the former case we are done. In the latter case, we conclude that the vast majority of nodes in $T_{b}$ have degree two.
@@
Let $G$ be an $n$-vertex triangulation of maximum-degree $\Delta$ and let $G^{*}$ be the dual of $G$. Then $G$ has a free set of size $\Omega(c(G^{*})/\Delta^{4})$
A result of Kant and Bodlaender [43] shows that one can add edges to any planar graph $G_{0}$ of maximum degree $\Delta$ to obtain a triangulation $G$ of maximum degree at most $\lceil 3\Delta/2\rceil+11$. Combined with 4, this establishes
@@
Let $G$ be an $n$-vertex planar graph of maximum-degree $\Delta$ and let $G^{*}$ be the dual of $G$. Then $G$ has a free set of size $\Omega(c(G^{*})/\Delta^{4})$.
2 shows that, for the class of planar graphs of maximum-degree $\Delta=O(1)$, the dual circumference and the size of the largest free set differ by only a constant factor. Combined with the best known lower bounds on the circumference of cubic triconnected graphs [44], 2 gives:
Corollary 4.
Let $G$ be an $n$-vertex triangulation of maximum-degree $\Delta$. Then $G$ has a free set of size $\Omega(n^{0.8}/\Delta^{4})$.

### v2

@@ 3.2 Free Sets in Planar Graphs
Therefore, $G_{x}:=G[\{x_{1},\ldots,x_{k}\}]$ is an induced outerplane subgraph of $G$ whose outer face is bounded by the cycle $C_{x}$. Each edge of $G$ with two endpoints in $G_{x}$ is contained in the closure of the interior of $C_{x}$. Thus each edge that is not in $E(G_{x})$ intersects the exterior of $C_{x}$ and has at most one endpoint on $C_{x}$. Let $E_{0}\subseteq E(G)\setminus E(G_{x})$ be the set of edges with no endpoint on $C_{x}$, and $E_{1}\subseteq E(G)\setminus E(G_{x})$ be the set of edges with exactly one endpoint on $C_{x}$. By Lemma 4, there is simple closed curve $C$ that contains a set $S$ of at least $k/2$ vertices of $G_{x}$, that is contained in the closure of the outer face of $G_{x}$, that contains $G_{x}$ in the closure of its interior, and that is proper and good with respect to $G_{x}$. These properties ensure that $C$ visits the vertices of $G_{x}$ in the same order as they appear on $C_{x}$ and that $C$ can be made to follow closely $C_{x}$ – in fact, so closely that $C$ intersects each edge of $E_{1}$ in one point and does not intersect any edge in $E_{0}$. Since the remaining edges of $G$ (those in $E(G_{x})$) are contained in the closure of the interior of $C_{x}$, the curve $C$ intersects each edge in $E(G)$ in at most one point, as illustrated in Fig. 8(d). Thus $C$ is a proper-good curve for $G$ that contains the vertices in $S$. By 1, $S$ is a free set in $G$ of size at least $k/2$.
@@
Next consider some maximal antichain $S:=y_{1},\ldots,y_{k}:=v_{i_{1}},\ldots,v_{i_{k}}$ of $(V(G),\prec)$ of length $k>1$ ordered by canonical ordering so that $i_{1}<\cdots<i_{k}$, as illustrated in Fig. 8(e). Since $k>1$, the set $S$ does not contain $v_{1}$ or $v_{2}$. Since $S$ is maximal and does not contain $v_{1}$ or $v_{2}$, we have $y_{k}=v_{n}$. Consider the sequence of cycles $C_{1},\ldots,C_{k}$ where $C_{j}$ is the cycle that bounds the outer face of $G_{i_{j}}$. By definition of canonical ordering, the interior of $C_{j}$ contains the interior of $C_{j-1}$ for each $j\in\{2,\ldots,k\}$. The nesting of these cycles is illustrated in Fig. 8(e), where each new colour shows the interior of $C_{j}$ that is not contained in the interior of $C_{j-1}$. Furthermore, the interior of $C_{j}$ must contain the vertex $y_{j-1}$ since, otherwise both $y_{j-1}$ and $y_{j}$ are on the outer face of $G_{i_{j}}$, which would mean that $y_{j-1}$ and $y_{j}$ are comparable. Consider now the union of two cycles $C_{j-1}$ and $C_{j}$ (or rather the union of the two closed curves that represent these two cycles in the crossing-free drawing of $G$). This union has one (bounded) face $f$ that contains both $y_{j-1}$ and $y_{j}$ on its boundary. The boundary of $f$ is a cycle $D_{j}$ in $G$ comprised of two paths: a path in $C_{j-1}$ containing $y_{j-1}$ and a path in $C_{j}$ containing $y_{j}$. Note that the union of all the cycles $D_{1},\dots D_{k}$, (or rather the union of the $k$ closed curves that represent these $k$ cycles in the crossing-free drawing of $G$) defines $k+1$ faces in the plane. These faces are illustrated in Fig. 8(f), where the interior of each face is assigned its own colour.
@@
We now construct a proper-good curve $C$ for $G$ that contains the vertices of $S$. Let $y_{0}$ be a point in the interior of the edge $v_{1}v_{2}$. Let $G_{1}^{\prime}:=G_{i_{1}}$ and, for each $j\in\{2,\ldots,k\}$, let $G_{j}^{\prime}$ be the near-triangulation whose outer face is bounded by the cycle $D_{j}$ described in the previous paragraph, and whose inner faces are faces of $G$. Graphs $G_{j}^{\prime}$ are illustrated in Fig. 8(f) by having their interiors shaded. Observe that $G^{\prime}_{1},\ldots,G^{\prime}_{k}$ have pairwise disjoint interiors. By Lemma 5, there is a simple curve $I_{j}$:$[0,1]\to\operatorname{int}(G^{\prime}_{j})$ with endpoints $y_{j-1}$ and $y_{j}$ that is proper with respect to $G^{\prime}_{j}$, for each $j\in\{1,\ldots,k\}$. Since $G^{\prime}_{1},\ldots,G^{\prime}_{k}$ have pairwise disjoint interiors, the curve obtained by taking the union of $I_{1},\ldots,I_{k}$ is simple and it is proper with respect to $G$. We complete this curve into a cycle by connecting $y_{k}$ to $y_{0}$ through the outer face of $G$. By 1, $S$ is a free set of size $k$.
@@
To complete the proof, we use Dilworth’s Theorem, which guarantees that the poset $(V(G),\prec)$ contains a chain of size at least $\sqrt{2n}$ or an antichain of size at least $\sqrt{n/2}$. In either case, we obtain a free set of size at least $\sqrt{n/2}$. ∎
@@ 3.3 Free Sets in Max-Degree-$\Delta$ Planar Graphs
Da Lozzo et al. [24] suggest the possibility that, since upper bounds on the circumference of dual graphs can be used to obtain upper bounds on the size of free sets, maybe lower bounds on circumference can be used to prove the existence of large collinear sets. Dujmović and Morin [27] show that, for planar graphs of bounded degree, this is indeed the case. In short, they show that a triangulation $G$ of maximum degree $\Delta$ whose dual has circumference $c(G^{*})$ has a free set of size $\Omega(c(G^{*})/\Delta^{4})$. A series of results has steadily improved the lower bounds on the circumference of $n$-vertex (not necessarily planar) cubic triconnected graphs [9, 14, 43, 12, 46]. The current record is held by Liu et al. [46] who show that, for any $n$-vertex cubic triconnected graph $G^{*}$, $c(G^{*})\in\Omega(n^{0.8})$.
In the remainder of this subsection, we describe some of the techniques used in [27] to establish the $\Omega(c(G^{*})/\Delta^{4})$ result. Cycles in $G^{*}$ are relevant to free sets because every cycle in $G^{*}$ corresponds to a proper-good curve in $G$. The resulting curve $C$ does not contain any vertices of $G$, but it is natural to try reroute $C$ to obtain a new curve that goes through some vertices of $G$. Fig. 9 (a) depicts a situation where this rerouting fails because the resulting curve is no longer proper because it intersects the red edge in Fig. 9(b) in two points. This leads to the following definition that lays out conditions under which such rerouting of curve $C$ is safe. We say that a vertex $v$ of $G$ is caressed by $C$ if the edges of $G$ that are incident to $v$ and intersected by $C$ appear consecutively around $v$, as depicted in Fig. 9(c) (the thick edges crossed by $C$ are consecutive around $v$ unlike in Fig. 9(a) where the set of edges crossed by $c$ is not consecutive around $v$). If $C$ caresses $v$ and intersects the edges $vv_{1},\ldots,vv_{r}$ (see Fig. 9(c)) then there is a sequence of faces $f_{0},\ldots,f_{r}$ where $vv_{i}$ is the edge shared between $f_{i-1}$ and $f_{i}$. Then the portion of $C$ that intersects $f_{0},\ldots,f_{r}$ can be replaced with a curve $C^{\prime}$ that enters $f_{0}$, proceeds directly to $v$, and immediately exits $f_{r}$ (see Fig. 9(d)). Because the original curve $C$ does not intersect any other edges incident to $v$, the modified curve $C^{\prime}$ is also a proper-good curve. This operation can be repeated on any set of caressed vertices that form an independent set:
@@
For each $v\in S$, reroute $C$ as described above so that $C$ contains $v$. Since $v$ is caressed by $C$, the rerouting that takes place at $v\in S$ causes $C$ to intersect each edge incident to $v$ in exactly one point, namely $v$, but does not change the intersection of $C$ with any edges not incident to $v$. Since the vertices in $S$ form an independent set, all of these rerouting operations do not cause $C$ to intersect any edge of $G$ in more than one point. Thus $C$ is a proper-good curve that contains $S$ so, by 1 $S$ is a free set in $G$. (The condition that $S$ has at least two vertices avoids the case in which $C$ crosses all the edges incident to a single vertex $v$.) ∎
The requirement that a curve $C$ caresses a vertex $v$ of $G$ is equivalent to requiring that the intersection of $C$ with the face $f_{v}$ of $G^{*}$ that contains $v$ is a path. When this happens, we say that $C$ caresses $f_{v}$. Thus finding a large collinear set in $G$ is equivalent to finding a cycle in $G^{*}$ that caresses many faces of $G^{*}$.
@@
Let $G$ be a triangulation and let $G^{*}$ be the dual of $G$. If some cycle $C$ in $G^{*}$ caresses $k$ faces of $G^{*}$, then $G$ contains a free set of size at least $k/4$.
@@
Unfortunately, there exist cubic triconnected planar graphs $G^{*}$ with no large faces that have a Hamiltonian cycle that caresses only four faces of $G^{*}$. One such graph is shown in Fig. 10. The main technical result of Dujmović and Morin [27] is to show that such cycles can be modified to produce cycles that caress many faces:
@@
Refer to Fig. 11. The proof of Lemma 8 is far too long to include here in any detail, so we give a high-level sketch. In the following informal sketch, we treat $\Delta$ as a fixed constant. Let $C$ be a cycle in $G^{*}$. Say that $C$ touches a face $f$ of $G^{*}$ if $C$ and $f$ share at least one edge. Since each face of $G^{*}$ has at most $\Delta$ edges and each edge of $C$ touches two faces of $G^{*}$ (one inside $C$ and one outside of $C$), the number of faces of $G^{*}$ that are touched by $C$ is at least $2|C|/\Delta$. At least $|C|/\Delta$ of these faces are in the interior of $C$ and at least $|C|/\Delta$ of these are in the exterior of $C$.
The proof defines a subgraph $H$ of $G^{*}$ that includes all the edges of $C$ and such that any cycle in the dual graph $H^{*}$ of $H$ contains faces inside and outside of $C$. Removing the edges from $H^{*}$ that correspond to edges of $C$ produces two trees $T_{0}$ and $T_{1}$, where $T_{0}$ contains faces of $H$ in the interior of $C$ and $T_{1}$ contains faces of $H$ in the exterior of $C$. For each $b\in\{0,1\}$, the tree $T_{b}$ has two important properties:
@@
The second of these properties says that any node of $T_{b}$ that contains many faces touched by $C$ either caresses many of these faces or has a high degree in $T_{b}$. The number of leaves in $T_{b}$ is $2+\sum_{x}(\deg_{T_{b}}(x)-2)$, where the sum is over all non-leaf nodes $x$ of $T_{b}$. Combining this with the first of these properties implies that $C$ caresses $\Omega(|C|)$ faces of $G^{*}$ or that $T_{b}$ has $\Omega(|C|)$ nodes but at most $\epsilon|C|$ leaves, for some small $\epsilon>0$. In the former case, we are done. In the latter case, we conclude that the vast majority of nodes in $T_{b}$ have degree two.
@@
Let $G$ be an $n$-vertex triangulation of maximum degree $\Delta$ and let $G^{*}$ be the dual of $G$. Then $G$ has a free set of size $\Omega(c(G^{*})/\Delta^{4})$.
2 shows that, for triangulations of maximum degree $\Delta=O(1)$, the dual circumference and the size of the largest free set differ by only a constant factor.
A result of Kant and Bodlaender [45] shows that one can add edges to any planar graph $G_{0}$ of maximum degree $\Delta$ to obtain a triangulation $G$ of maximum degree at most $\lceil 3\Delta/2\rceil+11$. Combining this with the best known lower bounds on the circumference of cubic triconnected graphs [46] and 2 gives:
@@
Let $G$ be an $n$-vertex planar graph of maximum degree $\Delta\geq 1$. Then $G$ has a free set of size $\Omega(n^{0.8}/\Delta^{4})$.